# SWE-CI: Evaluating Agent Capabilities in Maintaining Codebases via Continuous Integration

### 04 Mar 2026 v1

@@ Abstract
Large language model (LLM)-powered agents have demonstrated strong capabilities in automating software engineering tasks such as static bug fixing, as evidenced by benchmarks like SWE-bench. However, in the real world, the development of mature software is typically predicated on complex requirement changes and long-term feature iterations — a process that static, one-shot repair paradigms fail to capture. To bridge this gap, we propose SWE-CI, the first repository-level benchmark built upon the Continuous Integration loop, aiming to shift the evaluation paradigm for code generation from static, short-term functional correctness toward dynamic, long-term maintainability. The benchmark comprises 100 tasks, each corresponding on average to an evolution history spanning 233 days and 71 consecutive commits in a real-world code repository. SWE-CI requires agents to systematically resolve these tasks through dozens of rounds of analysis and coding iterations. SWE-CI provides valuable insights into how well agents can sustain code quality throughout long-term evolution.
@@ 1 Introduction
Building upon this insight, we propose SWE-CI (SoftWare Engineering – Continuous Integration), a novel benchmark designed to evaluate how well agents maintain code across long-term code evolution. SWE-CI comprises 100 tasks, each defined by a base commit and a target commit from a real-world repository, spanning on average 233 days and 71 consecutive commits of authentic evolutionary history. SWE-CI employs an Architect–Programmer dual-agent evaluation protocol: starting from the base commit, the agents execute a CI-loop that iteratively generates requirements, modifies source code, and runs tests, with the ultimate objective of passing all tests associated with the target commit. SWE-CI introduces EvoScore (Evolution Score) as a proxy metric: it measures functional correctness on future modifications, so that agents whose earlier decisions facilitate subsequent evolution score higher, while those that accumulate technical debt see progressively declining performance.
@@ 3.1 Data curation
Unlike SWE-Bench and similar benchmarks that draw exclusively from a handful of well-known open-source projects, we cast a wider net by searching across all Python repositories on GitHub. We then apply the following filtering criteria: (1) the repository has been actively maintained for at least three years; (2) it has accumulated more than 500 stars; (3) it contains configuration and dependency files (e.g., pyproject.toml and lockfiles) as well as a suite of unit tests; and (4) it is released under a permissive license such as MIT or Apache-2.0. After applying these filters, 4,923 repositories remain.
@@
The final SWE-CI benchmark comprises 100 samples drawn from 68 distinct repositories. On average, each base/oracle pair spans 233 days and 71 consecutive commits of real-world development history. In every pair, the transition from the base to the oracle codebase involves at least 500 lines of modified source code, excluding changes to test files. Each sample is shipped with the complete source code and a pre-built Docker environment to ensure reproducibility. These statistics confirm that SWE-CI captures substantial, long-term evolutionary episodes rather than trivial incremental changes.

### 17 Mar 2026 v2

@@ Abstract
Large language model (LLM)-powered agents have demonstrated strong capabilities in automating software engineering tasks such as static bug fixing, as evidenced by benchmarks like SWE-bench. However, in the real world, the development of mature software is typically predicated on complex requirement changes and long-term feature iterations — a process that static, one-shot repair paradigms fail to capture. To bridge this gap, we propose SWE-CI, the first repository-level benchmark built upon the Continuous Integration loop, aiming to shift the evaluation paradigm for code generation from static, short-term functional correctness toward dynamic, long-term maintainability. The benchmark comprises 100 tasks, each deriving from a real-world code repository with a development history spanning an average of 233 days and 71 consecutive commits. SWE-CI requires agents to systematically resolve these tasks through dozens of rounds of analysis and coding iterations. SWE-CI provides valuable insights into how well agents can sustain code quality throughout long-term evolution.
@@ 1 Introduction
Building upon this insight, we propose SWE-CI (SoftWare Engineering – Continuous Integration), a novel benchmark designed to evaluate how well agents maintain code across long-term code evolution. SWE-CI comprises 100 tasks, each defined by a base commit and a target commit from a real-world repository, with an average of 233 days and 71 consecutive commits of authentic evolutionary history in between. SWE-CI employs an Architect–Programmer dual-agent evaluation protocol: starting from the base commit, the agents execute a CI-loop that iteratively generates requirements, modifies source code, and runs tests, with the ultimate objective of passing all tests associated with the target commit. SWE-CI introduces EvoScore (Evolution Score) as a proxy metric: it measures functional correctness on future modifications, so that agents whose earlier decisions facilitate subsequent evolution score higher, while those that accumulate technical debt see progressively declining performance.
@@ 3.1 Data curation
Unlike SWE-Bench and similar benchmarks that draw exclusively from a handful of well-known open-source prjects, we cast a wider net by searching across all Python repositories on GitHub. We then apply the following filtering criteria: (1) the repository has been actively maintained for at least three years; (2) it has accumulated more than 500 stars; (3) it contains configuration and dependency files (e.g., pyproject.toml and lockfiles) as well as a suite of unit tests; and (4) it is released under a permissive license such as MIT or Apache-2.0. After applying these filters, 4,923 repositories remain.
@@
The final SWE-CI benchmark comprises 100 samples drawn from 68 distinct repositories. On average, each base/oracle pair is separated by 233 days and 71 consecutive commits of real-world development history. In every pair, the transition from the base to the oracle codebase involves at least 500 lines of modified source code, excluding changes to test files. Each sample is shipped with the complete source code and a pre-built Docker environment to ensure reproducibility. These statistics confirm that SWE-CI captures substantial, long-term evolutionary episodes rather than trivial incremental changes.
@@ 4.2 Results
Acknowledgements
The authors thank colleagues from Alibaba Group, including Jianan Ye, Cecilia Wang, Yijie Hu, Zongwen Shen, and Mingze Li, for their valuable suggestions and feedback on this paper.